# Video Instruction Tuning with Synthetic Data

### v2

@@ Abstract
footnotetext: $\heartsuit$ Work collaborated with ByteDance; $\P$ Co-senior authors
@@ 2 Related Work
The most related works to ours are the recent AI-generated synthetic video instruction tuning data, LLaVA-Hound (Zhang et al., 2024d) and ShareGPT4Video (Chen et al., 2024a), where they have used GPT-4 (OpenAI, 2023) to generate video captions and open-ended video question-answering. Although the quality of the captions and question-answer pairs has significantly improved, the video sources they use are too static to produce high-quality data for instruction-following scenarios. They also only use very sparse frames for prompting GPT-4V, which results in annotations that fail to capture nuanced actions and continuous plots in the videos. Additionally, Shot2Story (Han et al., 2023) and Vript (Han et al., 2023) also employ GPT-4V (OpenAI, 2023) for video captioning. Their outputs, however, include audio details, which are outside the scope of this study.
@@ 3.1 Video source
One important starting point in building a high-quality video instruction-following dataset is to find a sufficiently diverse pool of video data. From this pool, we can select the qualified videos. In our study of public video-language datasets—including video captioning, video question answering, video summarization, and moment-wise captioning—we noticed that although different datasets focus on various video understanding tasks (e.g., AGQA (Grunde-McLaughlin et al., 2021) for spatial-temporal relations and STAR (Wu et al., 2024a) for situational reasoning), most are sourced from ten main video sources. For instance, both AGQA and STAR use data from Charades (Sigurdsson et al., 2016). Specifically, these ten sources are HD-VILA-100M (Xue et al., 2022), InternVid-10M (Wang et al., 2023), VidOR (Shang et al., 2019), VIDAL (YouTube Shorts)(Zhu et al., 2023a), YouCook2(Zhou & Corso, 2017), Charades (Sigurdsson et al., 2016), ActivityNet (Caba Heilbron et al., 2015), Kinetics-700 (Kay et al., 2017), Something-Something v2 (Goyal et al., 2017), and Ego4d (Grauman et al., 2022). These sources offer a wide range of video data from different websites, viewpoints, and domains. The relationship between these ten selected video datasets and others is shown in Fig. 1. The videos from this ten datsets build the video pool for the further video selection. Notably, we use untrimmed videos from each source except for YouCook2 and Kinetics-700. We believe that cutting videos into clips can break the plot continuity, which is essential for understanding the videos.
@@
Figure 2: The video detail description creation pipeline. A three-level creation pipeline is considered, with each level developed via a recurrent approach. Note that $t$ is the index of time internal at its own level, and $T$ is the last time internal index. (a) To generate the caption for time internal $t$ at level-1, we condition on the current frames in this internal, the caption for time internal $t-1$, and the most recent description summary at level-2 if applicable. (b) To generate caption for time internal $t$ at level-2, we condtion on the previous caption at level-2, and captions from three most recent time internals at level-1. (c) To generate the overall caption at the last time internal $T$ at level-3, we condtion on the the most recent caption at level-2 and the current caption from level-1.
@@ 3.4 Dataset Statistics
We present the distribution in Figure 6. Our dataset shows a balanced mix across different video sources, providing a varied content selection. For each task type (caption, open-ended question, multiple-choice question), VIDAL (YouTube Shorts) has the highest share at 24.8%, 31.1%, and 30.1% respectively. It is followed by HD-VILA-100M (21.7%, 27.5%, 26.4%) and InternVid-10M (20.3%, 25.6%, 24.6%).
Figure 6 (Left) illustrates the distrubtion of the video duration. Video lengths range from 0s to 180s, with each length category containing at least 600 videos. Videos shorter than 50 seconds are numerous, mainly because all videos from VIDAL (24.8% of the dataset), which contains YouTube Shorts with lengths under 45 seconds. Figure 6 (Middle) illustrates the distribution on the number of words for the synthetic captions. Figure 6 (Right) shows how video length correlates with the length of captions. Generally, longer videos feature longer captions.
For each video in LLaVA-Video-178K, referencing InsTag (Lu et al., 2023), we employ an in-house tagging model to categorize the video content. Figure 7 displays the distribution of ten uniformly sampled video categories, showcasing examples from four of these categories. Among all videos, “comedy” predominates, primarily because YouTube Shorts is one of the most common sources in our dataset. Comedy is a typical genre that tends to attract high view counts—videos with large viewerships are more likely to be collected, as indicated in Table 1. Additionally, our dataset includes some domains less represented in current video-language datasets, such as computer games.
@@ 4 Experiments
For full evaluation, we consdier 11 video benchmarks. conducted tests across various video captioning , video open-ended question-answering and video multiple-choice question-answering benchmarks, including ActivityNet-QA (Yu et al., 2019), which features human-annotated action-related QA pairs from the ActivityNet dataset. We also utilized LongVideoBench (Wu et al., 2024b), EgoSchema (Mangalam et al., 2024), and MLVU (Zhou et al., 2024) for long video understanding, PerceptionTest (Pătrăucean et al., 2023) for assessing fine-grained perception skills, and VideoMME (Fu et al., 2024) and NExT-QA (Xiao et al., 2021) for diverse video domains and durations. Additional tests included VideoDetailCaption (LMMs-Lab, 2024) Dream-1K (Wang et al., 2024) for detailed video descriptions and Video-ChatGPT (Maaz et al., 2024) for visual chat.
For ablation studies in . 4.2 and Sec. 4.3, we conduct evaluation across 4 datasets. NExT-QA (Xiao et al., 2021) and PerceptionTest (Pătrăucean et al., 2023), which use training data from the LLaVA-Video-178K, are treated as in-domain datasets. Conversely, VideoMME (Fu et al., 2024) and EgoSchema (Mangalam et al., 2024) are considegreen zero-shot datasets.
@@ 4.1 Overall Results
test test test test test m-avg test mc val val wo/w-subs
@@
GPT-4V (OpenAI, 2023) 4.00 34.4 57.0 4.06 - 49.2 43.5 - - 61.3 59.9/63.3
GPT-4o (OpenAI, 2024) - 39.2 - - - 64.6 - - - 66.7 71.9/77.2
Gemini-1.5-Flash (Team et al., 2023) - 34.8 55.3 - 65.7 - - - - 61.6 70.3/75.0
Gemini-1.5-Pro (Team et al., 2023) - 36.2 57.5 - 72.2 - - - - 64.0 75.0/81.3
@@
VILA-40B (Lin et al., 2024) 3.37 33.2 58.0 3.36 58.0 - - 67.9 54.0 - 60.1/61.1
PLLaVA-34B (Xu et al., 2024a) - 28.2 60.9 3.48 - - 58.1 - - 53.2 -
LongVA-7B (Zhang et al., 2024c) 3.14 - 50.0 3.20 - 56.3 - 68.3 - - 52.6/54.3
IXC-2.5-7B (Zhang et al., 2024b) - - 52.8 3.46 - 37.3 69.1 71.0 34.4 - 55.8/58.8
LLaVA-OV-7B (Li et al., 2024c) 3.75 31.7 56.6 3.51 60.1 64.7 56.7 79.4* 57.1 56.5 58.2/61.5
VideoLLaMA2-72B (Cheng et al., 2024) - 27.1 55.2 3.16 63.9 61.2 62.0 - - - 61.4/63.1
LLaVA-OV-72B (Li et al., 2024c) 3.60 33.2 62.3 3.62 62.0 68.0 59.4 80.2* 66.9 61.3 66.2/69.5
LLaVA-Video-7B 3.66 32.5 56.5* 3.52 57.3 70.8 58.6 83.2* 67.9* 58.2 63.3/69.7
LLaVA-Video-72B 3.73 34.0 63.4* 3.62 65.6 74.4 64.1 85.4* 74.3* 61.9 70.5/76.9
@@ 4.2 Dataset Ablation
Note that the training set for LLaVA-Video includes six datasets: LLaVA-Video-178K, LLaVA-Hound (Zhang et al., 2024d), NExT-QA (Xiao et al., 2021), ActivityNet-QA (Yu et al., 2019), PerceptionTest (Pătrăucean et al., 2023), and image data from LLaVA-OneVision (Li et al., 2024c). In this section, we conduct ablation studies to assess the impact of each dataset. We separately fine-tune the LLaVA-OneVision (SI) model for each experimental setting, progressively adding datasets to the baseline. We use a video representation defined by $\mathcal{V}=(64,679,1,2)$
The results are presented in Table 3. Initially, we used a basic model trained solely on the LLaVA-Hound dataset as our baseline. Compared to this baseline, adding the LLaVA-Video-178K dataset significantly improved performance, enhancing scores in both in-domain and out-of-domain tasks. Specifically, we observed a 31.9-point increase in NExT-QA scores and a 9.1-point rise in VideoMME scores. Furthermore, including the PerceptionTest dataset significantly enhanced its associated task. Additionally, integrating high-quality image data provided modest benefits on EgoSchema.
@@
Method NExT-QA PerceptionTest EgoSchema VideoMME
@@ 4.3 Dataset Comparison
We conduct two ablation studies to further analyze our dataset and training strategy. As shown in Table 4, we compared three datasets where the language annotations are from GPT-4V/GPT-4o. For each experiment, we fine-tune the LLaVA-OneVision (SI) model separately on each specific dataset setting, utilizing a video representation defined by $\mathcal{V}=(64,679,1,2)$.
Two group of experiments are considered to assess the data quality of LLaVA-Video-178K compare to LLaVA-Hound and ShareGPT4Video. In the first group, to compare LLaVA-Video-178K with LLaVA-Hound, we randomly selected 900K open-ended questions to match the number in LLaVA-Hound. We included all captions and did not sample the multiple-choice questions. In the second group, comparing LLaVA-Video-178K to ShareGPT4Video, we randomly sampled 40K video captions to align with those in ShareGPT4Video. Since ShareGPT4Video lacks open-ended and multiple-choice questions, we supplemented with annotations from NExT-QA, PerceptionTest, and ActivityNet-QA. In the first group of Table 4, we compare LLaVA-Video-178K with LLaVA-Hound. Although LLaVA-Hound has more captions than LLaVA-Video-178K, our results are still better. As shown in Table 1, despite LLaVA-Hound annotates more videos, its quality is limited due to two main issues: (1) Static video: Its primary video source is WebVid (Bain et al., 2021), which tends to have relatively static content. (2) Sparse sampling: Although it includes data sources with dynamic videos, its sampling rate of 10 frames per video leads to annotations that do not fully capture the complete plot of the video. This underscores that the quality of video instruction-following data is more important than its quantity. Additionally, the second experiment group in Table 4 shows that the model trained with LLaVA-Video-178K outperforms that of ShareGPT4Video, highlighting the superiority of our data’s quality.
@@
#Caption #OE #MC NExT-QA PerceptionTest EgoSchema VideoMME
@@
LLaVA-Hound 900K 900k 0 64.4 51.4 51.0 51.0
LLaVA-Video-178K 178K 900k 0 73.2 (+8.8) 55.9 (+4.5) 49.8 (-1.2) 59.6 (+8.6)
@@ 5 Conclusion
This study introduces the LLaVA-Video-178K dataset, a high-quality synthetic dataset for video-language instruction-following. It is favored for its dense frame sampling rate in longer, untrimmed videos, covering diverse tasks such as captioning, open-ended and multi-choice QA. By training on the joint dataset of LLaVA-Video-178K with existing visual instruction tuning data, we developed a new model family, LLaVA-Video, which also considers video representation to effectively use GPU resources. This allows us to include more frames in the training process. The experimental results have demonstrated the effectiveness of the proposed synthetic dataset, and LLaVA-Video models have achieved excellent performance on a wide range of video benchmarks.
Appendix A Author Contributions
- Yuanhan Zhang contributed to the LLaVA-NeXT-Video series by developing the video training and inference codebase, creating the annotation pipeline, designing video representations, and conducting exploratory experiments.
- Jinming Wu contributed to the collecting and curating the video data, providing consistent technical support throughout the project, including helping improve codebase and conducting experiments.
- Wei Li engaged in designing the video data collection pipeline and experimental setups.
- Bo Li worked on providing LLaVA-OneVision codebase and training recipes as the initial foundation of current project, and offered help in integration to SGLang as well as demo service.
- Zejun Ma acquired a part of the data annotation and computational resources for the project.
- Ziwei Liu offered valuable suggestions throughout the project, contributing to discussions on the data collection pipeline design and experimental setups.
- Chunyuan Li initiated and led the project, designed the roadmap in the data collection, curation and expriments, and revised the paper.
Appendix B Video Representations
B.1 Efficient Video Representations in LMMs
@@
Appendix C Data
C.1 Video Detail Description
@@
C.2 Video Question Answering
@@ C.3 Dataset Comparison
Appendix D Beyond Singularity: Extensive Sampling Matters
@@ Appendix E Capabilities
Real-world Knowledge: Physical Laws [Uncaptioned image] User Please describe this video. LLaVA-Video The video features a person in a space station environment, wearing a dark gray shirt and a wristwatch. They are holding two containers filled with a thick, yellowish substance. The person tilts the containers to demonstrate the fluid’s behavior, showing how it stretches and moves between the containers. The background includes various equipment and panels typical of a space station, with labels such as ’GALLEY’ and ’FUSION’. The person continues to manipulate the containers, emphasizing the fluid’s properties. The video concludes with the person holding one container and twisting the lid to close it, further demonstrating the fluid’s properties. User Why doesn’t this yellowish substance fall? LLaVA-Video This yellowish substance doesn’t fall because it is in zero gravity.

### v3

# LLaVA-Video: Video Instruction Tuning With Synthetic Data

@@ 2 Related Work
The most related works to ours are the recent AI-generated synthetic video instruction tuning data, Islam et al. (2024) introduced Video ReCap, which recursively annotates video captions. Unlike Video ReCap, each clip-wise (level-1) description in our pipeline is generated with historical context. This ensures that connections from previous events in the video timeline are linked to the current event. LLaVA-Hound (Zhang et al., 2024d) and ShareGPT4Video (Chen et al., 2024a), where they have used GPT-4 (OpenAI, 2023) to generate video captions and open-ended video question-answering. Although the quality of the captions and question-answer pairs has significantly improved, the video sources they use are too static to produce high-quality data for instruction-following scenarios. They also only use very sparse frames for prompting GPT-4V, which results in annotations that fail to capture nuanced actions and continuous plots in the videos. Additionally, Shot2Story (Han et al., 2023) and Vript (Han et al., 2023) also employ GPT-4V (OpenAI, 2023) for video captioning. Their outputs, however, include audio details, which are outside the scope of this study.
@@ 3.1 Video source
One important starting point in building a high-quality video instruction-following dataset is to find a sufficiently diverse pool of video data. From this pool, we can select the qualified videos. In our study of public video-language datasets—including video captioning, video question answering, video summarization, and moment-wise captioning—we noticed that although different datasets focus on various video understanding tasks (e.g., , AGQA (Grunde-McLaughlin et al., 2021) for spatial-temporal relations and STAR (Wu et al., 2024a) for situational reasoning), most are sourced from ten main video sources. For instance, both AGQA and STAR use data from Charades (Sigurdsson et al., 2016). Specifically, these ten sources are HD-VILA-100M (Xue et al., 2022), InternVid-10M (Wang et al., 2023), VidOR (Shang et al., 2019), VIDAL (YouTube Shorts)(Zhu et al., 2023a), YouCook2(Zhou & Corso, 2017), Charades (Sigurdsson et al., 2016), ActivityNet (Caba Heilbron et al., 2015), Kinetics-700 (Kay et al., 2017), Something-Something v2 (Goyal et al., 2017), and Ego4d (Grauman et al., 2022). These sources offer a wide range of video data from different websites, viewpoints, and domains. The relationship between these ten selected video datasets and others is shown in Fig. 1. The videos from this ten datsets build the video pool for the further video selection. Notably, we use untrimmed videos from each source except for YouCook2 and Kinetics-700. We believe that cutting videos into clips can break the plot continuity, which is essential for understanding the videos.
@@
Figure 2: The video detail description creation pipeline. A three-level creation pipeline is considered, with each level developed via a recurrent approach. Note that $t$ is the index of time internal at its own level, and $T$ is the last time internal index. (a) To generate the caption for time internal $t$ at level-1, we condition on the current frames in this internal, the caption for time internal $t-1$, and the most recent description summary at level-2 if applicable. (b) To generate caption for time internal $t$ at level-2, we condition on the previous caption at level-2, and captions from three most recent time internals at level-1. (c) To generate the overall caption at the last time internal $T$ at level-3, we condition on the the most recent caption at level-2 and the current caption from level-1.
@@ 4 Experiments
For full evaluation, we consdier 11 video benchmarks. conducted tests across various video captioning , video open-ended question-answering and video multiple-choice question-answering benchmarks, including ActivityNet-QA (Yu et al., 2019), which features human-annotated action-related QA pairs from the ActivityNet dataset. We also utilized LongVideoBench (Wu et al., 2024b), EgoSchema (Mangalam et al., 2024), and MLVU (Zhou et al., 2024) for long video understanding, PerceptionTest (Pătrăucean et al., 2023) for assessing fine-grained perception skills, and VideoMME (Fu et al., 2024) and NExT-QA (Xiao et al., 2021) for diverse video domains and durations. Additional tests included VideoDetailCaption (LMMs-Lab, 2024), Dream-1K (Wang et al., 2024), Video-ChatGPT (Maaz et al., 2024) for detailed video descriptions, TemporalBench Cai et al. (2024) for fine-grained temporal understanding.
For ablation studies in . 4.2 and Sec. 4.3, we conduct evaluation across 4 datasets. NExT-QA (Xiao et al., 2021) and PerceptionTest (Pătrăucean et al., 2023), which use training data from the LLaVA-Video-178K, are treated as in-domain datasets. Conversely, VideoMME (Fu et al., 2024) and EgoSchema (Mangalam et al., 2024) are consider as zero-shot datasets.
@@ 4.1 Overall Results
test test test test test m-avg test mc val val m-acc wo/w-subs
@@
GPT-4o (OpenAI, 2024) - 39.2 - - - 64.6 - - - 66.7 35.3 71.9/77.2
Gemini-1.5-Pro (Team et al., 2023) - 36.2 57.5 - 72.2 - - - - 64.0 25.6 75.0/81.3
@@
VILA-40B (Lin et al., 2024) 3.37 33.2 58.0 3.36 58.0 - - 67.9 54.0 - - 60.1/61.1
PLLaVA-34B (Xu et al., 2024a) - 28.2 60.9 3.48 - - 58.1 - - 53.2 - -
LongVA-7B (Zhang et al., 2024c) 3.14 - 50.0 3.20 - 56.3 - 68.3 - - - 52.6/54.3
IXC-2.5-7B (Zhang et al., 2024b) - - 52.8 3.46 - 37.3 69.1 71.0 34.4 - 16.7 55.8/58.8
LLaVA-OV-7B (Li et al., 2024c) 3.75 31.7 56.6 3.51 60.1 64.7 56.7 79.4* 57.1 56.5 18.7 58.2/61.5
VideoLLaMA2-72B (Cheng et al., 2024) - 27.1 55.2 3.16 63.9 61.2 62.0 - - - - 61.4/63.1
LLaVA-OV-72B (Li et al., 2024c) 3.60 33.2 62.3 3.62 62.0 68.0 59.4 80.2* 66.9 61.3 26.6 66.2/69.5
LLaVA-Video-7B 3.66 32.5 56.5* 3.52 57.3 70.8 58.6 83.2* 67.9* 58.2 22.9 63.3/69.7
LLaVA-Video-72B 3.73 34.0 63.4* 3.62 65.6 74.4 64.1 85.4* 74.3* 61.9 33.7 70.5/76.9
@@ 4.2 Dataset Ablation
Note that the training set for LLaVA-Video includes six datasets: LLaVA-Video-178K, LLaVA-Hound (Zhang et al., 2024d), NExT-QA (Xiao et al., 2021), ActivityNet-QA (Yu et al., 2019), PerceptionTest (Pătrăucean et al., 2023), and image data from LLaVA-OneVision (Li et al., 2024c). In this section, we conduct ablation studies to assess the impact of each dataset. We separately fine-tune the LLaVA-OneVision (SI) model for each experimental setting, progressively adding datasets to the baseline.
The results are presented in Table 3. Initially, we used a basic model trained solely on the LLaVA-Hound dataset as our baseline. Compared to this baseline, adding the LLaVA-Video-178K dataset significantly improved performance, enhancing scores in both in-domain and out-of-domain tasks. Specifically, we observed a 31.9-point increase in NExT-QA scores and a 9.1-point rise in VideoMME scores. Furthermore, including the PerceptionTest dataset enhanced its associated task. Additionally, integrating high-quality image data provided modest benefits on EgoSchema.
@@
Method
NExT-QA
PercepTest
EgoSchema
VideoMME
@@ 4.3 Dataset Comparison
We conduct two ablation studies to analyze our dataset and training strategy. In Table 4, we compared three datasets where the language annotations are from GPT-4V/GPT-4o. For each experiment, we fine-tune the LLaVA-OneVision (SI) model separately on each specific dataset setting.
Two group of experiments are considered to assess the data quality of LLaVA-Video-178K compare to LLaVA-Hound and ShareGPT4Video. In the first group, to compare LLaVA-Video-178K with LLaVA-Hound, we randomly selected 900K open-ended questions to match the number in LLaVA-Hound. We included all captions and did not sample the multiple-choice questions. In the second group, comparing LLaVA-Video-178K to ShareGPT4Video, we randomly sampled 40K video captions to align with those in ShareGPT4Video. Since ShareGPT4Video lacks open-ended and multiple-choice questions, we supplemented with annotations from NExT-QA, PerceptionTest, and ActivityNet-QA. In the first group of Table 4, we compare LLaVA-Video-178K with LLaVA-Hound. Although LLaVA-Hound has more captions than LLaVA-Video-178K, our results are still better. The quality of LLaVA-Hound is limited due to two main issues: (1) Static video: Its primary video source is WebVid (Bain et al., 2021), which tends to have relatively static content. (2) Sparse sampling: its sampling rate of 10 frames per video leads to annotations that do not fully capture the complete plot of the video. This underscores that the quality of video instruction-following data is more important than its quantity. Additionally, the second experiment group in Table 4 shows that the model trained with LLaVA-Video-178K outperforms that of ShareGPT4Video, highlighting the superiority of our data’s quality.
@@
#Cap #OE #MC
NExT-QA
PercepTest
EgoSchema
VideoMME
@@
LLaVA-Hound 900K 900k 0 39.8 53.1 25.8 55.2
LLaVA-V-178K 178K 900k 0 73.2 55.9 49.8 59.6
@@ 5 Conclusion
This study introduces the LLaVA-Video-178K dataset, a high-quality synthetic dataset for video-language instruction-following. It is favored for its dense frame sampling rate in longer, untrimmed videos, covering diverse tasks such as captioning, open-ended and multi-choice QA. By training on the joint dataset of LLaVA-Video-178K with existing visual instruction tuning data, we developed a new model family, LLaVA-Video, which considers video representation to effectively use GPU resources. This allows us to include more frames in the training process. The experimental results have demonstrated the effectiveness of the proposed synthetic dataset, and LLaVA-Video models have achieved excellent performance on a wide range of video benchmarks.
6 Limitations
The videos in LLaVA-Video-178K are sourced from various platforms. This diversity introduces potential biases inherent in these sources. Furthermore, there is a concern regarding the potential skew in the question-answer pairs, possibly influenced by the annotators’ perspectives.
Acknowledgement
This study is supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOE-T2EP20221-0012, MOE-T2EP20223-0002), and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
Appendix A Video Representations
A.1 Efficient Video Representations in LMMs
@@
Appendix B Data
B.1 Video Detail Description
@@
B.2 Video Question Answering
@@ B.3 Dataset Comparison
Appendix C Beyond Singularity: Extensive Sampling Matters
@@ Appendix D Capabilities
Real-world Knowledge: Physical Laws
User Please describe this video.
LLaVA-Video The video features a person in a space station environment, wearing a dark gray shirt and a wristwatch. They are holding two containers filled with a thick, yellowish substance. The person tilts the containers to demonstrate the fluid’s behavior, showing how it stretches and moves between the containers. The background includes various equipment and panels typical of a space station, with labels such as ’GALLEY’ and ’FUSION’. The person continues to manipulate the containers, emphasizing the fluid’s properties. The video concludes with the person holding one container and twisting the lid to close it, further demonstrating the fluid’s properties.
User Why doesn’t this yellowish substance fall?
LLaVA-Video This yellowish substance doesn’t fall because it is in zero gravity.